In a certain community of 1,200 people, 60% are literate. Of the males, 50% are literates and of the females 70% are literate. What is the female population?
Answer by checkley77(12844)

.50x+.70(1,200-x)=.60*1,200
.50x+840x-.70x=720
-.20x=720-840
-.20x=-120
x=-120/-.20
x=600 males.
1,200-600=600 females.
Proof:
.50*600+.70*600=.60*1,200
300+420=720
720=720